# The long-term stability of the Vast Polar Structure and its connection to a possible previous passage of the LMC

### v1

@@ 3.4 Origin and evolution of the VPOS
Secondly, not only the LMC orbital path is well aligned with the VPOS, but also its passage by the MW does not disrupt or significantly alter the structure. The LMC is known to exert a very significant gravitational influence in the MW satellites, streams and the MW itself (see Vasiliev 2023 for a review). Although we observe some effects of such influence on the VPOS, namely the elongation of the distribution of MW satellites within the plane towards the LMC, a temporary compression, and a change in the orientation, the overall structural parameters of the VPOS remain largely unchanged. In other words, although the LMC has a substantial impact on the MW satellite system, it leaves the VPOS nearly intact. Moreover, once the LMC and its satellites are close to the MW they integrate into the plane without introducing drastic changes in its thickness, axis ratios or orientation. This smooth integration, combined with the mild effect of the arrival of the LMC suggests that the LMC system is not simply overlaid on the VPOS but may be fundamentally connected to its origin. Interestingly, it has been proposed in recent years that some of the brightest satellites of the MW may have been associated with the LMC in the past. For instance, studies based on orbital integration suggest that there is a small possibility that it could be the case of Carina and Fornax (Patel et al. 2020; Battaglia et al. 2022). However, Vasiliev (2024) highlights important limitations of orbit integration for tracing such associations, particularly since uncertainties in phase-space coordinates may bias against the association with the LMC. The spread that introduce the uncertainties is more likely to separate the galaxies from the LMC than associate them to it. Instead, Vasiliev (2024) employed N-body simulations of the interacting MW-LMC system to study the possible origins of observed MW satellites under the scenario where the LMC had performed two pericentres around the MW. The study shows that the probabilities of association with the LMC for the co-rotating on-plane MW satellites are not only non-negligible, but significant. In particular, for Carina, Draco, Fornax, and Ursa Minor, they are roughly 50%, i.e. they are as likely to have come with the LMC than to originate elsewhere. In the case of Crater II, Grus II and Tucana IV, the probability is $\sim 25\%$, lower but still significant. While not definitive proof of the VPOS’s origin, these findings are nonetheless compelling, particularly when viewed in the context of the broader body of evidence.
@@
Finally, we note that the thickness of the VPOS is very similar to the spatial extent of the LMC system. A simple estimate of this extent can be obtained by calculating the RMS or MAD of the distances of the LMC satellites to their host. However, this is a 3D measurement and therefore not directly comparable to the VPOS thickness, which is defined as the dispersion of projected distances of the galaxies to the plane. To enable a meaningful comparison, a consistent measurement must be applied to both systems. We measured the extent of the LMC system at $t=0$ using an iterative MC procedure consisting of $10^{3}$ realizations. In each iteration, the sky coordinates and distances of the LMC satellites were randomly sampled from normal distributions, and those sampled positions were then transformed into Galactocentric coordinates. To allow for a consistent comparison with the VPOS thickness, the 3D positions of the LMC satellites were projected onto a plane centered on the LMC, defined such that its normal vector aligns with the LMC’s velocity vector. We computed the RMS and MAD of the projected distances of the satellites to the LMC. This procedure was repeated in each iteration, and the final estimates were obtained by averaging the RMS and MAD values across all realizations. We find that the median spatial extent of the LMC group is $28.0^{+0.9}_{-1.0}$ kpc based on the RMS and $10.5^{+0.9}_{-0.7}$ kpc based on the MAD. It could be argued that the spatial extent of the LMC at $t=0$ may be not representative, since at present-time the LMC has just passed its pericentre and thus has been subjected to a very intense gravitational interaction recently. Therefore, we analogously calculated it for $t=-2$ Gyr, before it enters the MW halo. We measure similar values. For instance, under the potential L2M10 we obtain $35.9^{+4.4}_{+5.0}$ kpc and $9.6^{+3.6}_{-4.1}$ kpc measured for the RMS and MAD, respectively. The extent of the LMC system closely matches the thickness of the VPOS. While this may explain why the VPOS thickness remains largely unchanged when the LMC system approaches the MW, it also carries deeper implications. A key test of the group infall hypothesis is whether observed dwarf galaxy associations are compact enough to form thin planes (Pawlowski 2018). This has long been a challenge for the scenario, as most dwarf associations are too extended, typically spanning $\sim 150$–200 kpc, to produce thin structures like the VPOS (Metz et al. 2009). However, this is not the case for the LMC-system and the VPOS, where the spatial extent of both systems are clearly consistent. This suggests that the LMC system could be able to give rise to a thin structure, like the VPOS. Our results show that the LMC has not given rise to the VPOS during its recent passage, so considering all the evidence, it is possible that it was formed during a previous pericentre.
@@
To sum up, our study suggests a close connection between the VPOS and the LMC. We find that the VPOS has existed for at least the past 5 Gyr, well before the LMC crossed the MW’s virial radius roughly 2 Gyr ago. The alignment between the LMC’s orbital path and the VPOS plane, its endurance to the LMC gravitational influence, the smooth integration of LMC satellites into the structure during its approach, and the comparable spatial extents of both systems all point to a deeper physical association rather than a coincidental overlap. In the scenario where the LMC is currently having its second pericentric passage, and with that several of the on-plane MW satellites having a $\sim$50probability of past association with the LMC (Vasiliev 2024), it is plausible that the VPOS originated from satellites stripped from the LMC during its first pericentre. Such a scenario naturally explains the observed features, that otherwise appear as remarkable succession of very unlikely coincidences. However, while this scenario provides an elegant explanation for the origin of the VPOS, it remains possible that a number of the on-plane satellites joined the structure through other means. For example, the recent inclusion of Grus II in the VPOS, whose orbit appears to have been reoriented by the gravitational influence of the LMC, suggests that similar events could have occurred during a previous LMC pericentre, thereby aligning some additional MW satellites with the VPOS (a scenario analogous to the one explored in Garavito-Camargo et al. 2021). Another interesting case is that of Sculptor. In several models (V23, L3M10, and L3M11), Sculptor appears to have arrived from large distances and to have performed a single pericentre around the MW recently, yet is a clear member of the VPOS. While it cannot be ruled out that a limited number of satellites may align with the plane by chance, particularly given the size of the MW’s satellite population, Sculptor’s orbital history argues against a purely coincidental alignment. Our analysis shows that it came into close proximity with the LMC approximately $\sim$4 Gyr ago, within its tidal radius, suggesting that the LMC may have significantly influenced its current orbital configuration. Taken together, these considerations suggest that although the LMC may have primarily given rise to the VPOS during an early pericentre, some on-plane galaxies may be part of the VPOS due to a number of other phenomena, in which the LMC also seems to play a key role. We stress, however, that the origin of the VPOS is very complex question. In particular, the role of dwarf galaxy group accretion in formation planar satellite structures is still poorly understood, and deserves further investigation (Pawlowski, 2021).
3.5 Caveats
@@
We note that the primary source of systematic uncertainty in our analysis stems from our incomplete knowledge of the gravitational potentials of the MW and the LMC. Despite significant advances in the last years, these potentials remain poorly constrained (e.g. see Wang et al. 2020 for a review on the MW mass and potential). To address this, we employed six time-evolving potentials encompassing a realistic range of MW–LMC configurations based on current observational constraints. This allowed us to explore the stability and persistence of the VPOS under different MW-LMC configurations. In addition, we have repeated our analysis by exploring several variations of the V23 potential; using spherical and axisymmetric MW halos, as well as models with a lighter LMC ($5\times 10^{10}$ M) and triaxial and spherical MW DM halos. Across all models, we consistently find that the VPOS has remained a coherent and stable structure over at least the past 5 Gyr.
@@
4 Conclusions
In this work, we have studied the long-term stability and persistence of the Vast Polar Structure (VPOS). We started by evaluating the membership to the VPOS of 58 dwarf galaxies in the vicinity of the MW. Using accurate 6D phase-space data, we performed backward orbit integrations of the identified VPOS members under six different time-evolving gravitational potentials that account for the mutual interaction between the MW and the LMC. This allowed us to reconstruct the past trajectories of these galaxies over the last 5 Gyr in order to assess the evolution of the VPOS over time.
We identify 15 galaxies that are currently part of the VPOS according to their orbital poles, including 5 classical and 4 UFD satellites of the MW, as well as 6 likely members of the LMC satellite system. While the overall population of VPOS member galaxies appears stable over time, we find that Grus II has likely joined the structure only recently, as a consequence of the reorientation of its orbit induced by the gravitational influence of the LMC. Through the analysis of the structural properties of the VPOS over time, we find that:
  • The VPOS is a long-lived and stable structure. Over the last 5 Gyrs, the short-to-long axis ratio ($c/a\sim 0.2$), the thickness ($\sim$ 15 kpc), and the orientation of the VPOS remain roughly constant across all the tested potentials, and are consistent with previous observational estimates. This confirms that the VPOS is not a transient arrangement of some dwarfs around the MW, but a persistent and stable feature of the MW system.
  • The VPOS predates the LMC’s recent infall. Our analysis provides us with a lower limit for the age of the VPOS, it has existed for at least 5 Gyrs. Indeed, the MW satellites that currently form the VPOS, were already forming a planar structure well before the LMC crossed the MW’s virial radius. This rules out scenarios where the VPOS is a young phenomenon triggered by the recent passage of the LMC.
  • The approach of the LMC to the MW induces perturbations to the VPOS, but does not significantly alter it. While the LMC exerts a strong gravitational influence on the MW and its satellites during its infall, its impact on the VPOS is limited. We observe an elongation of the MW satellite distribution within the plane in the direction of the LMC, along with a temporary compression, and a shift in its orientation over the last Gyr. However, we note that these effects are subtle and do not significantly alter the structure or stability of the VPOS.
  • The structural properties of the VPOS remain consistent regardless of whether LMC satellites are included in the analysis. As the LMC and its satellites approach the MW, they blend smoothly with its on-plane satellite population, and the overall properties of the VPOS are mostly unchanged. At present, both configurations have very similar structural properties
In short, our results show that the VPOS is a stable, long-lived structure that predates the infall of the LMC and retains roughly constant structural properties over the last 5 Gyr.
Our findings provide compelling evidence about the possible origin of the VPOS, pointing to a close connection with the LMC and suggesting that this may not be its first passage by the MW. The good alignment between the LMC’s orbital path and the VPOS plane, as well as with the coherent motion of the MW satellites within it, would be an unlikely coincidence if the LMC were on its first approach. Yet, despite the LMC’s strong gravitational influence, the VPOS remains remarkably stable, and the arrival of LMC satellites does not alter its structure, further hinting at a causal connection. The close similarity between the spatial extent of the LMC satellite system and the VPOS thickness, would support that the LMC group could give rise to a structure like the VPOS, as a compact progenitor group is required to form such a thin plane. However, since our analysis shows that the VPOS predates the LMC’s recent infall, such a scenario would require the LMC to be at least on its second pericentric passage. Indeed, in such scenario, several MW satellites that are members of the VPOS have been shown to have a  50% probability of past association with the LMC (Vasiliev 2024), they are as likely of having come with the LMC than to form elsewhere. Therefore, the VPOS would be the result of the stripping of some of the LMC satellites during its first pericentre. This second-infall scenario naturally accounts for the observed alignment, limited perturbation by the LMC potential, and smooth integration between the MW on plane satellites and the LMC’s in the VPOS. However, while this scenario can plausibly explain the origin of the VPOS, some galaxies may have joined the structure in other ways. As seen in the case of Grus II, which likely joined the VPOS due to a recent reorientation of its orbit caused by the gravitational influence of the LMC, it cannot be ruled out that some satellites were similarly incorporated to the VPOS during an earlier pericentric passage of the LMC.
Although our results suggest to a strong connection between the VPOS and the LMC, we cannot directly trace the VPOS’s origin further back in time, as orbit integrations become unreliable beyond the timescales explored here. The role of group infall in assembling coherent planar structures like the VPOS remains largely unexplored, as does the detailed orbital history of the LMC system. Future Gaia data releases, with improved PM measurements, combined with cosmological simulations, will be key to understanding the origin of the VPOS and the broader context of the formation of planes of satellites in the LG and beyond.
Acknowledgements.
The authors thank E. Vasiliev for making publicly available the code AGAMA (Vasiliev 2018, 2019), and for his assistance with questions regarding its use. The authors acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S (MICIU/AEI/10.13039/501100011033), the Spanish Ministry of Science and Innovation project PID2021-124918NB-C41 (MICIU/AEI/10.13039/501100011033, FEDER, EU), and the RyC-MAX grant 20245MAX008 (CSIC). A. del Pino also acknowledges funding from the Ramón y Cajal fellowship RYC2022-038448-I (MICIU/AEI/10.13039/501100011033, co-funded by the European Social Fund Plus). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Appendix A The MW classic satellites and the stability of the VPOS

### v2

@@ 3.5 The LMC and the origin of the VPOS
Secondly, not only the LMC orbital path is well aligned with the VPOS, but also its passage by the MW does not disrupt or significantly alter the structure. The LMC is known to exert a very significant gravitational influence in the MW satellites, streams and the MW itself (see Vasiliev 2023 for a review). Although we observe some effects of such influence on the VPOS, the overall structural parameters remain largely unchanged. Moreover, once the LMC and its satellites are close to the MW they integrate into the plane without introducing drastic changes. All this suggests that the LMC system is not simply overlaid on the VPOS but may be fundamentally connected to its origin. Interestingly, it has been proposed in recent years that some of the brightest satellites of the MW may have been associated with the LMC in the past. For instance, studies based on orbital integration suggest that there is a small possibility that it could be the case of Carina and Fornax (Patel et al. 2020; Battaglia et al. 2022). However, Vasiliev (2024) highlights important limitations of orbit integration for tracing such associations, particularly since uncertainties in phase-space coordinates may bias against the association with the LMC. The spread that introduce the uncertainties is more likely to separate the galaxies from the LMC than associate them to it. Instead, Vasiliev (2024) employed N-body simulations of the interacting MW-LMC system to study the possible origins of observed MW satellites under the scenario where the LMC had performed two pericentres around the MW. The study shows that the probabilities of association with the LMC for the co-rotating on-plane MW satellites are not only non-negligible, but significant. In particular, for Carina, Draco, Fornax, and Ursa Minor, they are roughly 50%, i.e. they are as likely to have come with the LMC than to originate elsewhere. In the case of Crater II, Grus II and Tucana IV, the probability is $\sim 25\%$, lower, but still significant. While not definitive proof of the VPOS’s origin, these findings are nonetheless compelling, particularly when viewed in the context of the broader body of evidence.
@@
Finally, we note that the thickness of the VPOS is very similar to the spatial extent of the LMC system. We measured such extent at $t=0$ using an iterative MC procedure consisting of $10^{3}$ realizations. In each iteration, the sky coordinates and distances of the LMC satellites were randomly sampled from normal distributions, and then transformed into Galactocentric coordinates. To allow for a consistent comparison with the VPOS thickness, the 3D positions of the LMC satellites were projected onto a plane centered on the LMC, defined such that its normal vector aligns with the LMC’s velocity vector. We computed the RMS and MAD of the projected distances of the satellites to the LMC. This procedure was repeated in each iteration, and the final estimates were obtained by averaging the RMS and MAD values across all realizations. We find that the median spatial extent of the LMC group is $28.0^{+0.9}_{-1.0}$ kpc based on the RMS and $10.5^{+0.9}_{-0.7}$ kpc based on the MAD. It could be argued that the spatial extent of the LMC at $t=0$ may be not representative, since at present-time the LMC has just passed its pericentre and thus has been subjected to a very intense gravitational interaction recently. Therefore, we analogously calculated it for t=2t=-2 Gyr, before it enters the MW halo. We measure similar values. For instance, under the potential L2M10 we obtain $35.9^{+4.4}_{+5.0}$ kpc and $9.6^{+3.6}_{-4.1}$ kpc measured for the RMS and MAD, respectively. These values are well below the reported virial radius of the LMC ($\sim 75$150150 kpc; Kacharov et al. 2024) and are comparable to the typical median radial distance between LMC satellite analogues and their host found in cosmological simulations ($\sim 37$ kpc; Santos-Santos et al. 2021). The extent of the LMC system closely matches the thickness of the VPOS. While this may explain why the VPOS thickness remains largely unchanged when the LMC system approaches the MW, it also carries deeper implications. A key test of the group infall hypothesis is whether observed dwarf galaxy associations are compact enough to form thin planes (Pawlowski 2018). This has long been a challenge for the scenario, as most dwarf associations are too extended, typically spanning $\sim 150$–200 kpc, to produce thin structures like the VPOS (Metz et al. 2009). However, this is not the case for the LMC-system and the VPOS, where the spatial extent of both systems are clearly consistent. This suggests that the LMC system could be able to give rise to a thin structure, like the VPOS. Our results show that the LMC has not given rise to the VPOS during its recent passage, so considering all the evidence, it is possible that it was formed during a previous pericentre.
@@
In the scenario where the LMC is currently having its second pericentric passage, and with that several of the on-plane MW satellites having a $\sim$50% probability of past association with the LMC (Vasiliev 2024), it is plausible that the VPOS originated from satellites stripped from the LMC during its first pericentre. Such a scenario naturally explains the observed features (close alignment with the LMC orbit, mild influence, and smooth integration of LMC satellites), that otherwise appear as remarkable succession of very unlikely coincidences. However, while this scenario provides an elegant explanation for the origin of the VPOS, it remains possible that a number of the on-plane satellites joined the structure through other means. The recent inclusion of Grus II in the VPOS, whose orbit appears to have been reoriented by the gravitational influence of the LMC, suggests that similar events could have occurred during a previous LMC pericentre, thereby aligning some additional MW satellites with the VPOS (a scenario analogous to the one explored in Garavito-Camargo et al. 2021). Another interesting case is that of Sculptor. In several models (V23, L3M10, and L3M11), Sculptor appears to have arrived from large distances and to have performed a single pericentre around the MW recently, yet is a clear member of the VPOS. While it cannot be ruled out that a limited number of satellites may align with the plane by chance, particularly given the size of the MW’s satellite population, Sculptor’s orbital history argues against a purely coincidental alignment. Our analysis shows that it came into close proximity with the LMC approximately $\sim$4 Gyr ago, within its tidal radius, suggesting that the LMC may have significantly influenced its current orbital configuration. Taken together, these considerations suggest that although the LMC may have primarily given rise to the VPOS during an early pericentre, some on-plane galaxies may be part of the VPOS due to a number of other phenomena, in which the LMC also seems to play a key role. We stress, however, that the origin of the VPOS is very complex question. In particular, the role of dwarf galaxy group accretion in formation planar satellite structures is still not fully understood, and deserves further investigation (Pawlowski, 2021b).
4 Conclusions
In this work, we have studied the long-term stability and persistence of the Vast Polar Structure (VPOS). We started by evaluating the membership to the VPOS of 58 dwarf galaxies in the vicinity of the MW. Using accurate 6D phase-space data, we backward integrated the orbits of the identified VPOS members over the last 5 Gyrs under six different time-evolving gravitational potentials that account for the mutual interaction between the MW and the LMC. This allowed us to reconstruct the past trajectories of these galaxies in order to assess the evolution of the VPOS.
We identify 15 galaxies that are currently part of the VPOS, including 5 classical and 4 UFD satellites of the MW, as well as 6 likely members of the LMC satellite system. While the overall population of VPOS member galaxies appears stable over time, we find that Grus II has likely joined the structure only recently, as a consequence of the reorientation of its orbit induced by the gravitational influence of the LMC. Through the analysis of the structural properties of the VPOS over time, we find that:
  • The VPOS is not a transient arrangement of some dwarfs around the MW, but a long-lived and stable structure that has persisted over the last 5 Gyrs. The short-to-long axis ratio ($c/a\sim 0.2$), the thickness ($\sim$ 15 kpc), and the orientation of the VPOS remain roughly constant over time across all the tested potentials, and are consistent with previous observational estimates.
  • The VPOS predates the LMC’s recent infall. The MW satellites that are currently part of the VPOS, were already forming a planar structure well before the LMC crossed the MW’s virial radius. This rules out scenarios where the VPOS is a young phenomenon triggered by the recent passage of the LMC.
  • While the LMC infall exerts a strong gravitational influence on the MW and its satellites, it only induces mild perturbations in the VPOS that do not compromise the plane’s structure or long-term stability.
  • The structural properties of the VPOS remain consistent regardless of whether LMC satellites are included in the analysis. This indicates that, as the LMC group approached the MW, its satellites integrated smoothly with the existing MW on-plane population without significantly altering the plane’s overall configuration.
Our findings point to a close connection between the VPOS and the LMC, suggesting that the LMC’s current approach is not its first passage by the MW. The good alignment of the LMC’s orbit with the VPOS plane and the coherent motion of the MW satellites would be an unlikely coincidence if the LMC were on its first infall. Moreover, despite the LMC’s strong gravitational influence, the VPOS remains remarkably stable, and the arrival of LMC satellites leaves its structure largely unchanged, further hinting at a causal connection. The close similarity between the spatial extent of the LMC satellite system and the VPOS thickness supports the idea that it could have given rise to the VPOS, as forming such a thin plane requires a compact progenitor group. However, since our analysis shows that the VPOS predates the LMC’s recent infall, such a scenario would require the LMC to be at least on its second pericentric passage. Indeed, in such scenario, several MW satellites that are members of the VPOS have been shown to have a  50% probability of past association with the LMC (Vasiliev 2024), they are as likely of having come with the LMC than to form elsewhere. Therefore, the VPOS would be the result of the stripping of some of the LMC satellites during its first pericentre. This second-infall scenario naturally accounts for the observed alignment, limited perturbation by the LMC potential, and smooth integration between the MW on plane satellites and the LMC’s in the VPOS. However, while this scenario can plausibly explain the origin of the VPOS, some galaxies may have joined the structure in other ways. It is also possible that during an earlier passage, the LMC reoriented the orbits of some MW satellites, as appears to have occurred for Grus II recently.
Although our results suggest a strong connection between the VPOS and the LMC, we cannot directly trace the VPOS’s origin further back in time, as orbit integrations become unreliable beyond the timescales explored here. The role of group infall in assembling coherent planar structures like the VPOS remains an open question, as does the detailed orbital history of the LMC system. Future Gaia data releases, with improved PM measurements, combined with high-resolution cosmological simulations capable of resolving analogues of the LMC and the MW’s lowest-mass on-plane satellites, will be key to understanding the origin of the VPOS and the broader context of the formation of planes of satellites in the LG and beyond.
Acknowledgements.
The authors thank the anonymous referee for the comments that have helped to improve this paper. We thank E. Vasiliev for making publicly available the code AGAMA (Vasiliev 2018, 2019), and for his assistance with questions regarding its use. The authors acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S (MICIU/AEI/10.13039/501100011033), the Spanish Ministry of Science and Innovation project PID2021-124918NB-C41 (MICIU/AEI/10.13039/501100011033, FEDER, EU), and the RyC-MAX grant 20245MAX008 (CSIC). A. del Pino also acknowledges funding from the Ramón y Cajal fellowship RYC2022-038448-I (MICIU/AEI/10.13039/501100011033, co-funded by the European Social Fund Plus). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement.
Appendix A Caveats
@@
One intrinsic limitation of our study stems from the way we define VPOS membership: galaxies are classified based on the present-day direction of their orbital poles (see Section 2.2). This method inherently favors the selection systems that are already moving along a similar plane and whose uncertainties are small enough for their poles to remain within the VPOS area in the majority of the MC realizations. Such a selection naturally increases the likelihood of recovering a stable, long-lived planar structure. Additionally, as noted in Section 2.2, another limitation of this classification method is that it uses strict boundaries for the selection of on-plane galaxies, which can exclude potential VPOS members.
@@
In the orbit integration, satellites are treated as massless test particles in the time-dependent MW–LMC potential. This setup does not account for dynamical friction acting on the satellites themselves. While this omission could affect the reconstructed orbits of the most massive satellites (e.g. Fornax), we expect the impact on the VPOS analysis to be reduced, as the vast majority of satellites contributing to the structure are of sufficiently low mass that dynamical friction is negligible on the relevant timescales.
We note that the primary source of systematic uncertainty in our analysis stems from our incomplete knowledge of the gravitational potentials of the MW and the LMC. Despite significant advances in the last years, these potentials remain poorly constrained (e.g. see Wang et al. 2020 for a review on the MW mass and potential). To address this, we employed six time-evolving potentials encompassing a realistic range of MW–LMC configurations based on current observational constraints. This allowed us to explore the stability and persistence of the VPOS under different MW-LMC configurations. In addition to the potentials used in this work, we analysed the stability of the VPOS using some additional potentials based on V23. We first used a variation which keeps the same initial parametrization of the LMC but changes the DM halo of the MW (the baryonic component remains unchanged). In this case the MW halo follows truncated spherical NFW profile, with scale radius of 14 kpc and truncation radius of 300 kpc. We considered two additional potentials based in V23 and the spherical MW variation mentioned above, but considering a lighter LMC, with initial mass of $5\times 10^{10}$ M, scale radius of 5.61 kpc and truncation radius of 56.1 kpc (see Vasiliev 2023 for a more detailed explanation of these potentials). Under these additional potentials, we consistently find that the VPOS has remained a coherent and stable structure over at least the past 5 Gyr.
@@
Appendix B The MW classic satellites and the stability of the VPOS